OpenStudy (anonymous):

7 – 3x = 9 + x Please Help.

OpenStudy (lukecrayonz):

7-3x=9+x Okay:) So what is your first step? Or are you lost?

OpenStudy (anonymous):

Indeed I am. Lost, that is.

OpenStudy (lukecrayonz):

7-3x=9+x add 3x to the right side, by adding it to the left side (it's currently negative [subtraction]) 7=9+4x Subtract 9 from the right and to the left 7-9= -2 -2=4x Do you know how to solve from here?

OpenStudy (anonymous):

Like said, I need help. I can't quite grasp it.

OpenStudy (anonymous):

next you divide

OpenStudy (anonymous):

4x/4=x -2/4=-0.5 x=-0.5

OpenStudy (anonymous):

That should be your answer.

OpenStudy (anonymous):

You have to do the same thing to both sides of the =. You can do whatever you want (well, within the rules of mathematics). \[7 - 3x = 9 + x\]\[7 - 3x + 9000 = 9 + x + 9000\] That's valid. However, it doesn't help find x. To find x, we need to get it by itself. So do things that will help do that. \[7 - 3x = 9 + x\]\[7 - 3x + 3x = 9 + x + 3x\]\[7 = 9 + 4x\] That was actually helpful because it got rid of x's on one side.

OpenStudy (anonymous):

If I could, I'd give you all a Medal for taking your time to answer. Have a nice day. - Pdm